A REAR SIDE OF A TRUCK IS OPEN AND A BOX OF 40KG MASS IS PLACED 5M AWAY FROM THE OPEN END .THE COFFICIENT OF FRICTION BETWEEN THE BOX AND THE SURFACE BELOW IT IS 0.15.ON A STRAIGHT ROAD , A TRUCK STARTS FROM REST AND ACCELRATES WITH 2M/S2 . AT WHAT DISTANCE FROM THE STARTING POINT DOES THE BOX FALL OF THE TRUCK ? AFTER IGNORING THE MASS OF THE BOX

Grade:11

1 Answers

Aman
6 years ago
Take Right as + . Pseudo force on block=a trck×mas blck=80 normal on blck=wt=400 friction on blck =-60 net horizontal force =80-60=20 acceleration of block w r t trck=20/40=1/2 displacement of block w rt trck is 5 and initial velocity is 0.so time after blck fall= root 20 . S from start point=s of trck in time root20= 20